OpenStudy (anonymous):

Solve the equation: 8^x=20

OpenStudy (mimi_x3):

\[x = \frac{\ln20}{\ln8}=.. \]

OpenStudy (anonymous):

\[\ln (8^x)=\ln20\rightarrow x.\ln8=\ln20 \rightarrow x=(\ln20)/(\ln8)\]

OpenStudy (anonymous):

Do all of these answers get me the same thing??? Haha

OpenStudy (anonymous):

Also, I need help on 7^3x=18